OpenStudy (anonymous):

If one leg of a right triangle is 4 and the second leg is 6, then how long is the hypotenuse?

OpenStudy (anonymous):

((4^2+6^2))^(1/2)

OpenStudy (anonymous):

what?

OpenStudy (jennychan12):

in other words \[\sqrt{4^2+6^2}\]

OpenStudy (anonymous):

enter it on a calculator

OpenStudy (anonymous):

im not sure if I use the Pythagorean theorem or use 45 45 90?

OpenStudy (jennychan12):

x^1/2 = sqrt(x)

OpenStudy (jennychan12):

just do square root.

OpenStudy (anonymous):

and then I get the hypotenuse?

OpenStudy (jennychan12):

it's pythagorean. you're not given any other angles. yeah

OpenStudy (anonymous):

yuuup

OpenStudy (anonymous):

then would I get radical 52?

OpenStudy (anonymous):

but simplify it?

OpenStudy (jennychan12):

mhmm. then simplify that on your calcuator.

OpenStudy (anonymous):

look... c^2=a^2+b^2 where c is the hypotenuse, "a" and "b" are the "legs"

OpenStudy (anonymous):

I get that.

OpenStudy (anonymous):

so its 2 radical 13.

OpenStudy (jennychan12):

simplfity that as in decimals...

OpenStudy (anonymous):

7.211 units

OpenStudy (anonymous):

7.2

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thank you!!

OpenStudy (jennychan12):

no problem :)

OpenStudy (anonymous):

you're very welcome!